OpenStudy (anonymous):

1. What is the distance between A(6, 2) and B(9, 12)? Round to the nearest tenth. 2. What is the distance between M(–4, 2) and N(6, –3)? Round to the nearest tenth.

OpenStudy (michele_laino):

question #1 hint: distance d is: \[\large d = \sqrt {{{\left( {6 - 9} \right)}^2} + {{\left( {2 - 12} \right)}^2}} = ...?\]

OpenStudy (michele_laino):

since the general formula is: \[\large d = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} \]

OpenStudy (anonymous):

√109

OpenStudy (campbell_st):

do you know pythagoras' theorem... start by plotting the points  find the distances... and it avoids having to remember another formula